OpenStudy (anonymous):

What is the 32nd term of the arithmetic sequence where a1 = 13 and a13 = –59?

OpenStudy (campbell_st):

the term in an arithmetic sequence is \[T _{n}=a + (n -1)d\] us the information for the 13th term to find the common difference \[-59 = 13 + (13 -1 )d\] you need to solve -59 = 14 + 12d then with your value of d, a = 13 and n = 31, find the 32nd term by using the equation above.